From: Johnson Space Center
Posted: Monday, September 7, 2009
1. MPC Status - Due to the problems we had with the High Definition playbacks during last night's crew choice downlinks, we would like to perform the Digital Playback scheduled at MET 13:00 in Standard Definition. Following the ISS PAO event scheduled at MET 14:11, we will test the Shuttle MPC to ensure it is working properly before transfer to Station. Thus, pending the outcome of the troubleshooting, please hold off on transferring the MPC.
2. Shuttle Condensate Changeout - You will be changing out condensate CWC s/n 1070 with CWC s/n 5085 before hatch closure, but no earlier than 9/18:15. CWC s/n 5085 was previously used for EVA activities and should be transferred to the Middeck today during POST EVA XFER/RCNFG activities under MS2 and MS3. Stow the full CWC s/n 1070 at JPM1F7, on the rack front.
3. LiOH Changeout Cue Card - MSG 097 is your new LiOH Changeout Cue Card. Use this card starting on FD11. A new cue card is required due to changes made from the preflight plan.
4. CBM Status - Last night the CBM Prep for Demate was performed and everything checked out healthy and nominal. Since Bolt 4-1 is fully retracted from the interface after the R&R, we have Masked that bolt. Therefore, you will only need to perform your commanding verifications for only 15 of 16 bolts. Also, at the completion of the MPLM Demate, you may see a small amount of load (0.45kN) on Bolt 2-1. This is a known result of a drift in the load cell and is of no impact.
5. Replace Pages 3-114 Through 3-123.
// end //